# The Resolution of the Gauge Problem of Cosmology provides Insight into the Formation of the First Structures in the Universe

### v2

@@ VI.2 Linearized Equations
The system of linearized Einstein equations and conservation laws for closed, flat, and open flrw universes is succinctly expressed as follows:
@@
$\qquad\delta^{i}{}_{j}H\dot{h}^{k}{}_{k}+2\,{{}^{3}\!}R^{i}_{{\scriptscriptstyle(1)}j}=-\kappa\delta^{i}{}_{j}(\varepsilon_{\scriptscriptstyle(1)}-p_{\scriptscriptstyle(1)}),$ (12c)
@@
$\frac{1}{c}\frac{{\mathrm{d}}}{{\mathrm{d}}t}\left(\varepsilon_{\scriptscriptstyle(0)}(1+w)u^{i}_{\scriptscriptstyle(1)}\right)\,-$
@@
In these equations, the following notation is employed for the perturbed metric tensor: $h_{\mu\nu}:=-g_{{\scriptscriptstyle(1)}\mu\nu}$ and $h^{\mu\nu}:=g^{\mu\nu}_{{\scriptscriptstyle(1)}}$. In light of the use of synchronous coordinates, it follows that $h_{00}=0$ and $h_{0i}=0$. In addition, the raising and lowering of indices is performed by the spatial background metric tensor as defined in expression (5), namely, $h^{i}{}_{j}=g^{ik}_{\scriptscriptstyle(0)}h_{kj}$ where $g^{ik}_{\scriptscriptstyle(0)}=-\tilde{g}^{ik}/a^{2}$. A vertical bar denotes covariant differentiation with respect to the spatial background metric tensor $g_{{\scriptscriptstyle(0)}ij}=-a^{2}\tilde{g}_{ij}$ or, equivalently, with respect to $\tilde{g}_{ij}$. In the context of flrw universes, where $\Gamma^{k}_{{\scriptscriptstyle(0)}ij}=\tilde{\Gamma}^{k}{}_{ij}$, the operations of taking the time derivative and the covariant derivative are shown to commute. Equation (12a) represents the linearized energy density constraint equation (linearized Friedmann equation), while Eqs. (12b) correspond to the linearized momentum constraint equations. The linearized dynamical equations are presented in Eqs. (12c). Equations (12d) and (12e) represent the linearized energy density conservation law and the linearized energy-momentum conservation laws, respectively. Finally, the linearized particle number density conservation law is given by Eq. (12f).
As previously stated in Sec. II, the pressure perturbation, the spatial curvature perturbation, and the covariant divergence of the spatial fluid velocity play a pivotal role in a cosmological perturbation theory. In consequence, the expressions for these quantities are now derived.
@@ VI.2.1 Pressure Perturbation
The gauge-dependent perturbation to the pressure (8) is given by
@@
The quantity representing the true physical perturbation, designated as $p^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}$, will be derived in Sec. XI on thermodynamics.
@@ VI.2.2 Spatial Curvature Perturbation
Lifshitz’ expression for the perturbed connection coefficients presented in Ref. [2], Eq. (I.3) and Ref. [31], Eq. (10.9.1) is a tensor, which is given by
@@
By combining (14) and (15), the following expression for the perturbed spatial Ricci tensor may be derived:
@@
By raising the index $i$, one arrives at the following result, employing (11b):
@@
By employing $g_{\scriptscriptstyle(0)}^{ij}h^{k}{}_{i\mid j\mid k}=g_{\scriptscriptstyle(0)}^{ij}h^{k}{}_{i\mid k\mid j}$, one arrives at the contraction of the perturbed spatial Ricci tensor:
@@
In Sec. VII it will be demonstrated that the expression (18) represents the local perturbation to the spatial curvature ${{}^{3}\!}R_{{\scriptscriptstyle(0)}}$, or the spatial curvature perturbation, induced by local density perturbations. The spatial curvature perturbation ${{}^{3}\!}R_{\scriptscriptstyle(1)}$ is incorporated into the evolution equations (60), as shown in Eq. (128).
VI.2.3 Covariant Divergence of the Spatial Fluid Velocity
In the background flrw universe, the fluid four-velocity is given by $u^{\mu}_{\scriptscriptstyle(0)}=\delta^{\mu}{}_{0}$, so that the perturbation $\theta_{\scriptscriptstyle(1)}$ to the expansion scalar $\theta:=u^{\mu}{}_{;\mu}$ is
@@ VI.2.3 Covariant Divergence of the Spatial Fluid Velocity
where $\vartheta_{\scriptscriptstyle(1)}$ is the covariant divergence of the spatial part $u^{i}_{\scriptscriptstyle(1)}$ of the perturbed fluid four-velocity $u^{\mu}_{\scriptscriptstyle(1)}$. In deriving (19), the following identity is employed: $(u^{k}{}_{|k})_{{\scriptscriptstyle(1)}}=u^{k}_{{\scriptscriptstyle(1)}|k}$, which follows from $u^{i}_{{\scriptscriptstyle(0)}}=0$. The covariant divergence $\vartheta_{{\scriptscriptstyle(1)}}$ is incorporated into the evolution equations (60), as shown in Eq. (133).
VII Decomposition of Spatial Tensors
The following important theorem is used to decompose the linearized Einstein equations and conservation laws into three independent systems. In Refs. [34, 36, 3], it is demonstrated that any symmetric spatial tensor ${\cal{T}}_{ij}$ of rank two can be unambiguously decomposed in a covariant manner into three irreducible components:
@@ VII Decomposition of Spatial Tensors
The constituents possess the following properties:
@@
where $\phi(t,\bm{x})$ and $\zeta(t,\bm{x})$ are two independent potentials. Consequently, the component $h^{i}_{\parallel\,j}$ can be expressed in terms of the two independent potentials as follows:
$h^{i}_{\parallel\,j}=\frac{2}{c^{2}}\bigl{(}\phi\delta^{i}{}_{j}+\zeta^{\mid i}{}_{\mid j}\bigr{)}.$ (22)
The factor $2/c^{2}$ is included in view of the nonrelativistic limit. Evidently, from (21), it can be inferred that an expression analogous to (22) exists for the tensor ${{}^{3}\!}R^{i}_{{\scriptscriptstyle(1)}\parallel\,j}$. In the context of cosmological perturbation theory, only the contraction ${{}^{3}\!}R_{{\scriptscriptstyle(1)}\parallel}$ of the latter tensor is required. Upon substituting (22) into (18), the following result is obtained:
@@
$\theta_{{\scriptscriptstyle(1)}}=\vartheta_{{\scriptscriptstyle(1)}}-\dfrac{1}{c^{2}}\bigl{(}3\dot{\phi}+\dot{\zeta}^{|k}{}_{|k}\bigr{)}.$ (24)
As evidenced by (VII), the perturbed Ricci scalar ${{}^{3}\!}R_{{\scriptscriptstyle(1)}\parallel}$ encompasses two distinct potentials $\phi(t,\bm{x})$ and $\zeta(t,\bm{x})$. In the nonrelativistic limit, as will be discussed in detail in Sec. IX, the potential $\phi$ becomes independent of time and equal to the Newtonian potential, while the potential $\zeta$ becomes inconsequential.
The spatial component $\bm{u}_{\scriptscriptstyle(1)}$ of the perturbed fluid four-velocity can be decomposed as follows
@@
~$\bm{\tilde{\nabla}}\bm{\cdot}\bm{u}_{\scriptscriptstyle(1)}=\bm{\tilde{\nabla}}\bm{\cdot}\bm{u}_{{\scriptscriptstyle(1)}\parallel},\quad\bm{\tilde{\nabla}}\bm{\times}\bm{u}_{\scriptscriptstyle(1)}=\bm{\tilde{\nabla}}\bm{\times}\bm{u}_{{\scriptscriptstyle(1)}\perp}.$ (26)
In these expressions, the generalized vector differential operator, denoted by $\bm{\tilde{\nabla}}$, is defined by the relation ~${\tilde{\nabla}_{i}v_{k}:=v_{k\mid i}}$. The expressions (25)–(26) are referred to as the Helmholtz decomposition. The following section will demonstrate that the decomposition of ${{}^{3}\!}R^{i}_{{\scriptscriptstyle(1)}j}$ and $h^{i}{}_{j}$ as given by (20) and the properties (21) are consistent with the the Riemann tensor (11a) and the momentum constraint equations (12b).
It is important to note that the velocity vector $\bm{u}_{{\scriptscriptstyle(1)}\parallel}$ is not uniquely determined by the system of equations (12). Since $\bm{u}_{{\scriptscriptstyle(1)}\parallel}$ is irrotational, it is possible to supplement it with the gradient of an arbitrary function. Given that the system of equations (12) is invariant under the gauge transformation (1), where the gauge function $\xi^{\mu}(t,\bm{x})$ is given by (6), it follows that
@@
is also a solution of the system (12). As a result, the gauge modes associated with $\bm{u}_{{\scriptscriptstyle(1)}\parallel}$ are given by
@@
This result is of particular significance in the context of the derivation of the nonrelativistic limit, as presented in Sec. IX.
VII.1 Decomposition of the Linearized Equations
The system of equations (12) is split into three independent systems by using the decomposition (20) with the properties (21) for $h_{ij}$ and ${{}^{3}\!}R_{ij}$, as well as (25)–(26). The solutions to these systems are conventionally designated as tensor perturbations $\ast$, vector perturbations $\perp$, and scalar perturbations $\parallel$.
@@ VII.1 Decomposition of the Linearized Equations
VII.1.1 Tensor Perturbations
From the expressions (18) and (21), it can be deduced that ${{}^{3}\!}R_{{\scriptscriptstyle(1)}\ast}=0$. This, in combination with Eq. (12a), yields ${\varepsilon_{{\scriptscriptstyle(1)}}=0}$. Eqs. (12b) yield the result that ${\bm{u}_{\scriptscriptstyle(1)}=\bm{0}}$. This, in turn, implies, with (19) that $\theta_{{\scriptscriptstyle(1)}}=0$. From (4a), it follows that ${\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}=0}$. Given that $\theta_{{\scriptscriptstyle(1)}}=0$, Eqs. (10d) and (12f) are identical. This implies that $n_{{\scriptscriptstyle(1)}}=0$. This, in turn, implies that $n^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}=0$. Finally, the result obtained from either (12c) or (12d) is that $p_{\scriptscriptstyle(1)}=0$. In consequence, the evolution equations for tensor perturbations are as follows:
@@ VII.1.1 Tensor Perturbations
VII.1.2 Vector Perturbations
From the expressions (18) and (21), it can be deduced that ${{}^{3}\!}R_{{\scriptscriptstyle(1)}\perp}=0$. This implies that
$h^{kl}_{\perp\mid k\mid l}=0.$ (30)
Moreover, it is found from Eq. (12a), that ${\varepsilon_{{\scriptscriptstyle(1)}}=0}$. Raising the index $i$ of Eqs. (12b) with $g^{ij}_{\scriptscriptstyle(0)}$, and subsequently taking the covariant derivative with respect to the index $j$, one finds, using that $\dot{g}^{ij}_{\scriptscriptstyle(0)}=-2Hg^{ij}_{\scriptscriptstyle(0)}$,
$\dot{h}^{kj}_{\perp{\mid k\mid j}}+2Hh^{kj}_{\perp{\mid k\mid j}}=-2\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)u^{j}_{{\scriptscriptstyle(1)}\mid j}.$ (31)
The combination of (30) and (31) results in ~$\bm{\tilde{\nabla}}\bm{\cdot}\bm{u}_{\scriptscriptstyle(1)}=0$. This indicates that only the component $\bm{u}_{{\scriptscriptstyle(1)}\perp}$ remains. From (19) one finds that $\theta_{{\scriptscriptstyle(1)}}=0$. This, combined with $\varepsilon_{{\scriptscriptstyle(1)}}=0$, yields $\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}=0$. From $\theta_{{\scriptscriptstyle(1)}}=0$ one finds that Eq. (12f) is identical to the background equation (10d), implying that $n_{{\scriptscriptstyle(1)}}=0$. This, in turn, implies with (4a) that $n^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}=0$. Finally, the result obtained from either (12c) or (12d) is that $p_{\scriptscriptstyle(1)}=0$. Therefore, the system of equations for vector perturbations is given by
@@ VII.1.2 Vector Perturbations
$\!\ddot{h}^{i}_{\perp j}+3H\dot{h}^{i}_{\perp j}+2R^{i}_{{\scriptscriptstyle(1)}\perp j}=0,$ (32b)
@@
VII.1.3 Scalar Perturbations
In this particular instance one has ${{}^{3}\!}R_{{\scriptscriptstyle(1)}\parallel}\neq 0$. By taking the covariant derivative of (12b) with respect to the index $j$ and subsequently substituting the expression (22), the following result is obtained:
@@ VII.1.3 Scalar Perturbations
Upon interchanging the indices $i$ and $j$ and subtracting the result from Eqs. (33) one finds
@@
where it is used that $\dot{\phi}_{\mid i\mid j}=\dot{\phi}_{\mid j\mid i}$ and $\dot{\zeta}^{\mid k}{}_{\mid k\mid i\mid j}=\dot{\zeta}^{\mid k}{}_{\mid k\mid j\mid i}$. By rearranging the covariant derivatives, Eqs. (34) can be transformed in the following form:
$(\dot{\zeta}^{\mid k}{}_{\mid i\mid k\mid j}-\dot{\zeta}^{\mid k}{}_{\mid i\mid j\mid k})-(\dot{\zeta}^{\mid k}{}_{\mid j\mid k\mid i}-\dot{\zeta}^{\mid k}{}_{\mid j\mid i\mid k})\,+$
@@
The expressions for the commutator of second-order covariant derivatives (see Ref. [31], Sec. 6.5) are given by
@@
Upon substituting the background Riemann tensor (11a), it is found that the left-hand sides of Eqs. (35) vanished identically. This implies that ~×$\bm{\tilde{\nabla}}\bm{\times}\bm{u}_{\scriptscriptstyle(1)}=\bm{0}$, so that only the component $\bm{u}_{{\scriptscriptstyle(1)}\parallel}$ remains. Since ~$\bm{\tilde{\nabla}}\bm{\cdot}\bm{u}_{{\scriptscriptstyle(1)}\parallel}\neq 0$ implies that $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}\neq 0$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}\neq 0$, it can thus be concluded that only scalar perturbations are coupled to density perturbations.
VIII Evolution Equations for Scalar Perturbations
As demonstrated in the preceding section, the evolution of $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ is determined by the background equations (10) and the equations that govern the evolution of scalar perturbations. In this section, the findings of the preceding section are utilized to reformulate the system (12) into a novel set of equations that exclusively describe the evolution of scalar perturbations. Given the subsequent focus on scalar perturbations, the subscript $\parallel$ is omitted. The evolution equations for scalar perturbations are as follows:

### v3

@@ VI.2 Linearized Equations
The system of linearized Einstein equations and conservation laws for closed, flat, and open flrw universes is given by
@@
[…]
@@
[…]
@@
In these equations, the perturbed metric tensor is given by $h_{\mu\nu}:=-g_{{\scriptscriptstyle(1)}\mu\nu}$ and $h^{\mu\nu}:=g^{\mu\nu}_{{\scriptscriptstyle(1)}}$. In light of the use of synchronous coordinates, it follows that $h_{00}=0$ and $h_{0i}=0$. Raising and lowering of indices is performed by the spatial background metric tensor as defined in expression (8), namely, $h^{i}{}_{j}=g^{ik}_{\scriptscriptstyle(0)}h_{kj}$ where $g^{ik}_{\scriptscriptstyle(0)}=-\tilde{g}^{ik}/a^{2}$. A vertical bar denotes covariant differentiation with respect to the spatial background metric tensor $g_{{\scriptscriptstyle(0)}ij}=-a^{2}\tilde{g}_{ij}$. In the context of flrw universes, where $\Gamma^{k}_{{\scriptscriptstyle(0)}ij}=\tilde{\Gamma}^{k}{}_{ij}$, the operations of taking the time derivative and the covariant derivative are shown to commute. Equation (13a) represents the linearized energy density constraint equation (linearized Friedmann equation), while Eq. (13b) corresponds to the linearized momentum constraint equation. The linearized dynamical equations are presented in Eqs. (13c). Equations (13d) and (13e) represent the linearized energy density conservation law and the linearized momentum conservation law, respectively. Finally, the linearized particle number density conservation law is given by Eq. (13f).
Pressure perturbations, spatial curvature perturbations, and the covariant divergence of the fluid velocity play pivotal roles in the evolution of density perturbations. Therefore, these quantities must be included in the perturbation theory. The next three subsections will derive expressions for these quantities.
@@ VI.2.1 Pressure Perturbation
The gauge-dependent quantity $p_{{\scriptscriptstyle(1)}}$, which is associated with pressure (2), is given by
@@
The quantity $p^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}$ representing the pressure perturbation will be derived in Sec. XII on thermodynamics.
@@ VI.2.2 Spatial Curvature Perturbation
Lifshitz’s expression for the perturbed connection coefficient presented in Ref. [2], Eq. (I.3) and Ref. [42], Eq. (10.9.1) is a tensor, which is given by
@@
By combining (15) and (16), the following expression for the perturbed spatial Ricci tensor can be derived (see Ref. [42], Eq. (10.9.3):
@@
By raising the index $i$ and using (12b), one arrives at
@@
By using $g_{\scriptscriptstyle(0)}^{ij}h^{k}{}_{i\mid j\mid k}=g_{\scriptscriptstyle(0)}^{ij}h^{k}{}_{i\mid k\mid j}$, one arrives at the contraction of the perturbed spatial Ricci tensor:
@@
In Sec. VII it will be demonstrated that the expression (19) represents the local perturbation to the spatial curvature ${{}^{3}\!}R_{{\scriptscriptstyle(0)}}$ induced by local density perturbations. The quantity ${{}^{3}\!}R_{\scriptscriptstyle(1)}$ will be used in Sec. IX to eliminate the metric perturbations and their spatial derivatives from the momentum constraint equation (13b). It is incorporated into the perturbation theory using Eq. (127).
VI.2.3 Covariant Divergence of the Fluid Velocity
[…]
@@ VI.2.3 Covariant Divergence of the Fluid Velocity
where $\vartheta_{\scriptscriptstyle(1)}$ is the covariant divergence of the fluid vector. The quantity $\theta_{{\scriptscriptstyle(1)}}$ will be used in Sec. IX to eliminate the time derivative of the metric perturbation from the energy density constraint equation (13a) in favor of $\vartheta_{{\scriptscriptstyle(1)}}$. The latter quantity is incorporated into the perturbation theory using Eq. (132).
VII Decomposition of Spatial Tensors and Vectors
The system of linearized Einstein equations and conservation laws (13) is not suitable for deriving the perturbation theory because it contains solutions related to and unrelated to density perturbations. Therefore, our goal is to transform the system (13) into one that is exclusively linked to density perturbations and excludes all other solutions. To that end, we will use two key theorems to break the linearized equations down into three independent systems. These theorems are presented in the next two subsections.
VII.1 Tensor Decomposition
Refs. [45, 46, 47] demonstrate that any rank-two symmetric spatial tensor ${\cal{T}}_{ij}$ can be decomposed into three irreducible components in a covariant manner:
@@ VII.1 Tensor Decomposition
The components have the following properties:
@@
where $\phi(t,\bm{x})$ and $\zeta(t,\bm{x})$ are two independent potentials. Consequently, the component $h^{i}_{\parallel\,j}$ of the perturbed metric can be expressed in terms of the two independent potentials as follows:
[…]
The factor $2/c^{2}$ is included in view of the Newtonian limit. By substituting (23) into (19), the following result is obtained:
@@
[…]
The expression (VII.1) reveals that two potentials, $\phi(t,\bm{x})$ and $\zeta(t,\bm{x})$, are associated with cosmological density perturbations in open and closed flrw universes. For a flat flrw universe, however, density perturbations are characterized by one potential, $\phi(t,\bm{x})$. This is because, for scalar perturbations, the quantity $\theta_{{\scriptscriptstyle(1)}}$ can be eliminated from expressions (6a) and Eqs. (37), as was done in (122) and (123).
As will be demonstrated in Sec. X, in the Newtonian limit, the potential $\phi(t,\bm{x})$ becomes identical to the Newtonian potential, $\phi(\bm{x})$.
VII.2 Vector Decomposition
Since $\bm{u}_{\scriptscriptstyle(1)}$ is a vector in three-space, it can be decomposed using the Helmholtz decomposition as follows:
@@ VII.2 Vector Decomposition
$\bm{\tilde{\nabla}}\bm{\cdot}\bm{u}_{\scriptscriptstyle(1)}=\bm{\tilde{\nabla}}\bm{\cdot}\bm{u}_{{\scriptscriptstyle(1)}\parallel},\quad\bm{\tilde{\nabla}}\bm{\times}\bm{u}_{\scriptscriptstyle(1)}=\bm{\tilde{\nabla}}\bm{\times}\bm{u}_{{\scriptscriptstyle(1)}\perp}.$ (27)
[…]
Since $\bm{u}_{{\scriptscriptstyle(1)}\parallel}$ is irrotational, it can be supplemented with the gradient of an arbitrary function. Given that the system of equations (13) is invariant under the gauge transformation (3), where the gauge function $\xi^{\mu}(t,\bm{x})$ is given by (9), it follows that
@@
is also a solution to the system (13). As a result, the gauge modes associated with $\bm{u}_{{\scriptscriptstyle(1)}\parallel}$ are given by
@@
This result will be used in Sec. X on the Newtonian limit.
VIII Decomposition of the Linearized Equations
The system of equations (13) is split into three independent systems by using the decomposition (21) with the properties (22) for $h_{ij}$ and ${{}^{3}\!}R_{ij}$, as well as the decomposition (26) with the properties (27) for $\bm{u}_{{\scriptscriptstyle(1)}}$. The solutions to these systems are conventionally designated as tensor perturbations $\ast$, vector perturbations $\perp$, and scalar perturbations $\parallel$.
@@ VIII Decomposition of the Linearized Equations
VIII.1 Tensor Perturbations
From the expressions (19) and (22), it follows that ${{}^{3}\!}R_{{\scriptscriptstyle(1)}\ast}=0$, and Eq. (13b) yields $\bm{u}_{\scriptscriptstyle(1)}=\bm{0}$}, so that, with (20), $\theta_{{\scriptscriptstyle(1)}}=0$. Equation (13a) yields $\varepsilon_{{\scriptscriptstyle(1)}}=0$}. From (6a), it follows that $\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}=0$}. Given that $\theta_{{\scriptscriptstyle(1)}}=0$, Eqs. (11d) and (13f) are identical, so $n_{{\scriptscriptstyle(1)}}=0$. This, in turn, implies with (6a), that $n^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}=0$. Finally, the result obtained from either (13d) or the contraction of (13c) is that $p_{{\scriptscriptstyle(1)}}=0$. Consequently, the evolution equations for tensor perturbations are as follows:
@@ VIII.1 Tensor Perturbations
VIII.2 Vector Perturbations
Expressions (19) and (22), imply that, for vector perturbations, we have ${{}^{3}\!}R_{{\scriptscriptstyle(1)}\perp}=-h^{kj}_{\perp|k|j}$. Raising the index $i$ of the momentum constraint equation (13b) with $g^{ij}_{\scriptscriptstyle(0)}$, and subsequently taking the covariant derivative with respect to the index $j$, one finds, using that $\dot{g}^{ij}_{\scriptscriptstyle(0)}=-2Hg^{ij}_{\scriptscriptstyle(0)}$,
$\dot{h}^{kj}_{\perp{\mid k\mid j}}+2Hh^{kj}_{\perp{\mid k\mid j}}=-2\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)u^{j}_{{\scriptscriptstyle(1)}\perp\mid j}.$ (31)
The right-hand side of this equation vanishes, as can be seen from (27). Consequently, the covariant divergence of the momentum constraint equation (31) can only be satisfied if ${{}^{3}\!}R_{{\scriptscriptstyle(1)}\perp}=0$. The consequences are as follows.
Equation (13a) implies that $\varepsilon_{{\scriptscriptstyle(1)}}=0$}, and (20) yields $\theta_{{\scriptscriptstyle(1)}}=0$. The latter identity shows that Eq. (13f) is identical to the background equation (11d), so $n_{{\scriptscriptstyle(1)}}=0$}. From (6a) we find that $\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}=0$ and $n^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}=0$. Finally, either (13d) or the contraction of (13c) implies that $p_{{\scriptscriptstyle(1)}}=0$. Therefore, the system of equations for vector perturbations is given by
@@ VIII.2 Vector Perturbations
$\!\ddot{h}^{i}_{\perp j}+3H\dot{h}^{i}_{\perp j}+2\,{{}^{3}\!}R^{i}_{{\scriptscriptstyle(1)}\perp j}=0,$ (32b)
@@
VIII.3 Scalar Perturbations
In this particular instance one has ${{}^{3}\!}R_{{\scriptscriptstyle(1)}\parallel}\neq 0$, as follows from expression (VII.1). By taking the covariant derivative of the momentum constraint equation (13b) with respect to the index $j$ and subsequently substituting the expression (23), the following result is obtained:
@@ VIII.3 Scalar Perturbations
By interchanging the indices $i$ and $j$ and subtracting the result from Eqs. (33) one finds
@@
where it is used that $\dot{\phi}_{\mid i\mid j}=\dot{\phi}_{\mid j\mid i}$ and $\dot{\zeta}^{\mid k}{}_{\mid k\mid i\mid j}=\dot{\zeta}^{\mid k}{}_{\mid k\mid j\mid i}$. Rearranging the second-order covariant derivatives converts Eq. (34) into a form with only commutators on the left-hand side:
[…]
@@
The commutators of second-order covariant derivatives (see Ref. [42], Sec. 6.5) are given by
@@
Upon substituting the background Riemann tensor (12a), the left-hand side of Eq. (35) vanishes identically. This implies that $\bm{\tilde{\nabla}}\bm{\times}\bm{u}_{\scriptscriptstyle(1)}=\bm{0}$, so that only the component $\bm{u}_{{\scriptscriptstyle(1)}\parallel}$ remains. Since $\bm{\tilde{\nabla}}\bm{\cdot}\bm{u}_{{\scriptscriptstyle(1)}\parallel}\neq 0$ implies that $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}\neq 0$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}\neq 0$, it can thus be concluded that only scalar perturbations are coupled to density perturbations.
IX Evolution Equations for Scalar Perturbations
As demonstrated in the preceding section, the evolution of $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ is determined by the background equations (11) and the equations that describe the evolution of scalar perturbations. In this section, the findings from the previous section are used to reformulate the system (13) into a new system of equations that exclusively describes the evolution of scalar perturbations. Given the subsequent focus on density perturbations, the subscript $\parallel$ is omitted. The evolution equations for scalar perturbations are as follows: